The difference between the ice point and steam point in a liquid in glass thermometer is 30cm. What temperature is recorded when the mercury thread is 10cm above the ice point?


1
Expert's answer
2022-05-17T08:57:32-0400
"t=\rm 10\: cm\times \frac{100^\circ C-0^\circ C}{30\: cm}=33^\circ C"